41) The risk-free return is 5.9% and the market risk premium is 10.4%. What is the expected return for the following portfolio? (State your answer in percent with two decimal places.)
Stock      Beta      Investment
AAA        3.8       $500,000
BBB        2.3       $1,100,000
CCC        1.8       $1,500,000
DDD        1.0       $1,000,000

20.62%
26.52%
14.82%
8.92%
19.83%
Answer #1

Total investment =(500,000+1,100,000+1,500,000+1,000,000)=$4,100,000

Portfolio beta=Respective beta*Respective investment weight

=(500,000/4,100,000*3.8)+(1,100,000/4,100,000*2.3)+(1,500,000/4,100,000*1.8)+(1,000,000/4,100,000*1)

=1.982926829

required return= risk-free rate +Beta*market risk premium

=5.9+1.982926829*10.4

=26.52%(Approx).